Solve this system of equations using the elimination method:
3a-3b=-15
-3a-3b=-3

Found 2 solutions by mananth, Devonte12:
Answer by mananth(16949)

3a-3b=-15
-3a-3b=-3
..
add up the equations
3a-3a-3b-3b=-15-3
-6b=-18
/-6
b=-18/-6
b=3
..
plug value of b in first equation
3a-3b=-15
3a-3*3=-15
3a-9=-15
+9
3a=-15+9
3a=-6
/3
a=-2
(-2,3)
..
m.ananth@hotmail.ca

Answer by Devonte12(1)

r=-3s
r+4s=10
plug in -3s into r
so you get -3sr + 4s =10
add r's up and you get 1s=10
s=10
same for next problem
x-y=5
x+2y=7
isolate one of the variables
-y=5-x
divide by -
so you get y=-5+x
plug in for y
and solve

p l+l+w+w or 2l+2w
l= w+24
so plug in to ^^
2(w +24) +2w=268
so you get
4w +48= 268
subtract by 48 and divide by 4
and you find w
plug in w into l+w+24
sorry but i didnt want to do them all if you really need help with them ill do it but next time try to do less